How do you slove a equation with varibles on both sides?
Example:
1/2n+16=2/3n-4

Answer by stanbon(75887)

(1/2)n+16=(2/3)n-4
Get the variables on one side and the numbers on the other:
Subtract 16 from both sides to get:
(1/2)n=(2/3)n-4-16
Combine the -4 and the -16
(1/2)n=(2/3)n-20
Subtract (2/3)n from both sides to get:
(1/2)n-(2/3)n=-20
[1/2 - 2/3]n = -20
[(3/6)-(4/6)]n=-20
(-2/3)n=-20
Multiply both sides by -3/2 to get:
n=60
---------
Cheers,
Stan H.